STRUCTURE AND SPECTROSCOPIC CONSTANTS OF X̃^{2}Π NCS: AN AB INITIO MOLECULAR OR-BITAL STUDY

\underline{TSUNEO HIRANO} and UMPEI NAGASHIMA, Grid Technology Research Center, Institute of Advanced Industrial Science and Technology, 6-9-3 Ueno, Taito-ku, Tokyo 110-0015, Japan; TINA ERICA ODAKA and PER JENSEN, FB 9 – Theoretische Chemie, Bergische Universität – Gesamthochschule Wuppertal, D-42097 Wuppertal, Germany.

Much is not known for NCS, while its sister molecule NCO is a quite familiar molecule for high resolution molecular spectroscopy and is known as one of the typical molecules to show the Renner effect. Molecular constants of X̃^{2}Π NCS are known,^{a} but there has been a discrepancy between theoretically predicted and experimentally determined rotational constants.^{b} At the MR-SDCI+Q and MR-ACPF levels with the full-valence active space, our calculated B_{0} values were smaller by at least 0.4 % than experimentally observed B_{0} value, as is the case reported by Ouazbir, et al.^{b} When the core-valence correlation is included, this situation was improved to give an error of 0.05% in B_{e} and B_{0}. The predicted B_{e} and B_{0}, though preliminary at the moment, at the level of core-valence MR-SDCI+Q/aug-cc-pCVQZ are 6113.9 and 6103.5 MHz, respectively, against the experimental B_{0}value^{a} of 6106.62162(25) MHz. The N-C and C-S equilibrium bond lengths are predicted to be 1.178 and 1.632 Å, respectively, and hence the B_{0} of the isotopemer NC^{34}S could be 5959 MHz.

aT. Amano and T. Amano, J. Chem. Phys., 95, 2275-2279 (1991); and references therein .

bM. Ouazbir, G. Chambaud, P. Rosmus and P.J. Knowles, Phys. Chem. Chem. Phys., 1, 2649-2655 (1999).